1/2

Heinrich Heine,  Heinrich Heine,  Heinrich Heine,
The Lorelei  The Lorelei  Die Lorelei
tradukita de Mark Twain  tradukita de L. W. Garnham

An ancient legend of the Rhine

I cannot divine what it mea-

neth, This haunting nameless pain: A tale of the bygone ages Keeps brooding through my

brain:

I do not know what it signifies. That I am so sorrowful? A fable of old Times so terrifies, Leaves my heart so thoughtful.

Ich weiß nicht, was soll es be-

deuten, daß ich so traurig bin; ein Märchen aus alten Zeiten, das kommt mir nicht aus dem

Sinn.

The faint air cools in the glo-

aming, And peaceful flows the Rhine, The thirsty summits are drin-

king The sunset’s flooding wine; The air is cool and it darkens, And calmly flows the Rhine; The summit of the mountain

hearkens In evening sunshine line.

Die Luft ist kühl und es dun-

kelt, und ruhig fließt der Rhein; der Gipfel des Berges funkelt im Abendsonnenschein.

The loveliest maiden is sitting High-throned in yon blue air, Her golden jewels are shining, She combs her golden hair; The most beautiful Maiden ent-

rances Above wonderfully there, Her beautiful golden attire

glances, She combs her golden hair.

Die schönste Jungfrau sitzet dort oben wunderbar, ihr goldnes Geschmeide blit-

zet, sie kämmt ihr goldenes Haar.

She combs with comb that is

golden, And sings a weird refrain That steeps in a deadly en-

chantment The listener’s ravished brain:

With golden comb so lustrous, And thereby a song sings, It has a tone so wondrous, That powerful melody rings.

Sie kämmt es mit goldenem

Kamme, und singt ein Lied dabei; das hat eine wundersame, gewaltige Melodei.

The doomed in his drifting

shallop, Is tranced with the sad sweet

tone, He sees not the yawing brea-

kers, He sees but the maid alone:

The shipper in the little ship It effects with woe sad might; He does not see the rocky slip, He only regards dreaded height.

Den Schiffer im kleinen Schiffe ergreift es mit wildem Weh; er schaut nicht die Felsenriffe, er schaut nur hinauf in die

Höh.

…  …  …

2/2

The pitiless billwos engulf

him!- So perish sailor and bark; And this, with her baleful sin-

ging, Is the Lorelei’s gruesome work.

I believe the turbulent waves Swallow the last shipper and

boat; She with her singing craves All to visit her magic moat.

Ich glaube, die Wellen ver-

schlingen am Ende Schiffer und Kahn; und das hat mit ihrem Singen die Lorelei getan.

Traduko de la Germana poemo “Die Lorelei” de Heinrich Heine (∗1797-12-13 – †1856-02-17) en la Anglan de Mark Twain.

Arg-2-11 (2003-10-13 04:42:59)

A Tramp Abroad. Vol 1-2. Leibzig: Tauchnitz, 1880 Band I, Mark Twain 1880 (rf. http: // www. loreley. com/ loreley/ marctwai. htm )

Traduko de la Germana poemo “Die Lorelei” de Heinrich Heine (∗1797-12-13 – †1856-02-17) en la Anglan de L. W. Garnham.

Arg-2-14 (2003-10-11 23:04:57)

L.W. Garnham, Bachelor of Arts, LEGENDS OF THE RHINE; mentioned in: A Tramp Abroad. Vol 1-2. Leibzig: Tauchnitz, 1880 Band I, Mark Twain 1880 Verkinto de tiu ĉi Germana poemo estas Heinrich Heine (∗1797-12-13 – †1856-02-17), publikigita 1823.

Arg-2-2 (2014-04-23 18:09:02)

Aus  dem  Werk  “Buch

der Lieder”, Kapitel “Die Heimkehr”, Abschnitt II. Siehe auch ein http: // gutenberg. spiegel. de unter: http: // gutenberg. spiegel. de/ ?id= 5&x d= 1131&kapitel= 12&cHash= eb14eca2272# gb_found . Vidu ankaŭ: http: // www. gedichte. vu/ ?die_lorelei. html . La melodion verkis la komponisto Friedrich Silcher. Vidu la retejon http: // www. lieder-archiv. de/ die_lorelei-notenblatt_ 300493. html .

http://www.poezio.net/version?poem-id=2&version-id=11,14,2